OpenStudy (anonymous):

Elizabeth has a hat filled with three colored beads green, purple, and pink. Elizabeth randomly pulls tiles out of the bag , records the color , and places the tile back into the bag. She repeats this 50 times and recorded the results in the table above. What is the experimental probability that Elizabeth will pull out a purple tile?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

green 18 purple 26 pink 6

OpenStudy (anonymous):

first you got add all the tiles up 18+26+6=50

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

dont you divide by 3 after you add them all up

OpenStudy (anonymous):

no

OpenStudy (igreen):

Where's the table? @Marc1313

OpenStudy (anonymous):

and since she pulled 50 tiles and the # of tiles oare fifty you wouldn`t even have to change the bottom # on the fraction. it would be purple tiles over / 50

OpenStudy (anonymous):

green 16 purple 26 pink 6 @iGreen

OpenStudy (anonymous):

it would be purple tiles over / 50

OpenStudy (igreen):

Okay, are you sure? Because that's only 48 times..

OpenStudy (anonymous):

that is 50 times ?? @iGreen

OpenStudy (anonymous):

no 18+26+6= 50

OpenStudy (igreen):

No..16 + 26 + 6 = 48 Check the table again..

OpenStudy (igreen):

Oh, you made a mistake..first you said 18 then you said 16..

OpenStudy (anonymous):

oh sorry its 18 plus 26 plus 6

OpenStudy (igreen):

Yeah..so just write the number of purple tiles divided by the total number of tiles(50).

OpenStudy (anonymous):

so 50

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

the answer is 52%

OpenStudy (igreen):

Yep

OpenStudy (anonymous):

you divide the 26/50

OpenStudy (igreen):

26/50 = 0.52 = 52%

OpenStudy (anonymous):

yay!

OpenStudy (igreen):

Nice work \(\huge\color{lime}\checkmark\)